OpenStudy (anonymous):

Put the following radical expression into simplified form. Assume the variable represents a positive number.

OpenStudy (anonymous):

\[\frac{ 4\sqrt{12x^4} }{ \sqrt{3x} }\] @whpalmer4 , can you tell me if I'm doingit right?

OpenStudy (anonymous):

doing it*

OpenStudy (anonymous):

\[\frac{ 4(2x^2 \sqrt{3}) }{ \sqrt{3x} }*\frac{ \sqrt{3x} }{ \sqrt{3x} }\]

OpenStudy (anonymous):

\[\frac{ 8x^2\sqrt{6x} }{ 3x }\]?

OpenStudy (anonymous):

That's far as I got, I'm not sure if it's suppose to be 6x

OpenStudy (whpalmer4):

The first step is correct, the second is not.

OpenStudy (whpalmer4):

\[4(2x^2\sqrt{3})*\sqrt{3x} = 4(2x^2\sqrt{3})*\sqrt{3}*\sqrt{x}\]does that give you a hint as to the right direction?

OpenStudy (anonymous):

I just know the x loses its radical. Was the top part correct? the 8x^2?

OpenStudy (whpalmer4):

that last line I gave you is the numerator expression — can you simplify it for me?

OpenStudy (anonymous):

\[8x^2\] I got lost after this...

OpenStudy (anonymous):

@Jack1

OpenStudy (anonymous):

.